Calculate (3t) * t^4 where * denotes convolution.

(3t)*t^4 =
Answer #1

elo convolution o (3+)* ť = 24 (3(+-2)) da ( assuming) te [0,oo) = 23t24 -25] da = ft 3 xt -as] da 3 [t-? 10 17 = 3 t 3 30 (3